OpenStudy (fools101):

I need help!!! Change 16°C to the nearest degree Fahrenheit. can anyone help me with this please! thank u

OpenStudy (stormfire1):

the formula is F = (C * 9/5) + 32 so it will be F = (16 * 9/5) + 32 which = 60.8 degrees F or 61 degrees (nearest degree)